# Collimation of Fast Radio Burster 20201124A; Repeaters vs. Apparent Non-Repeaters

### v2

@@ 4 Discussion
Alignment of the magnetic dipole and rotation axes of repeating FRB would explain the difficulty of finding their rotational periodicity because radiation from a field structure azimuthally symmetric around the rotation axis would not be rotationally modulated. It is implausible that neutron star fields are accurately dipolar near the surface, so this argument suggests that FRB are emitted at many neutron star radii. Emission along open field lines is consistent with, but does not require, that the periodic activity of some repeating FRB results from free precession of the neutron star (Zanazzi & Lai, 2020).
@@
The phenomenological differences (Pleunis et al., 2021) between bursts from repeaters and apparent non-repeaters, such as the “sad trombone” effect, must then be attributed to differences in the electrodynamics of aligned and mis-aligned rotators. This is consistent with the assumption that FRB are powered, like pulsars, by spin-down energy, not magnetostatic energy.

### v3

@@ 4 Discussion
Alignment of the magnetic dipole and rotation axes of active repeating FRB, as previously suggested (Luo et al., 2025), would explain the difficulty of finding their rotational periodicity because radiation from a field accurately azimuthally symmetric around the rotation axis would not be rotationally modulated. It is implausible that neutron star fields are accurately dipolar near the surface, so this argument suggests that FRB are emitted at many neutron star radii. Emission along open field lines is consistent with, but does not require, that the periodic activity modulation of some repeating FRB results from free precession of the neutron star (Zanazzi & Lai, 2020).
@@
The phenomenological differences (Pleunis et al., 2021) between bursts from repeaters and from apparent non-repeaters, such as the “sad trombone” effect, may then be attributed to differences in the electrodynamics of aligned and mis-aligned rotators. This is consistent with the assumption that FRB are powered, like pulsars, by spin-down energy, not magnetostatic energy.